# The connection of polaritonic chemistry with the physics of a spin glass

### v1

@@ 4.2 Probing the existence of a polarization glass phase
4.2.1 Equilibrium Susceptibilities
The hallmark of replica symmetry breaking in spin glass theory can be attributed to the emergence of two different static equilibrium susceptibilities, 78 which are observed in experiments 90 and in the SK model. 30 The two extreme cases describe the response of the system subject to a small external field perturbation. In the so-called zero-field cooled case, the system remains inside a given state while changing the magnetization (electric) field, with a corresponding linear response susceptibility $\chi_{LR}$. In contrast, the (true) thermodynamic equilibrium susceptibility $\chi_{eq}$ describes the situation, where the spin glass is allowed to relax to the thermodynamically most favored state in presence of a weak external field perturbation.
@@ 4.2.1 Equilibrium Susceptibilities
Experimentally, the static linear-response susceptibility $\chi_{LR}$ can be measured by looking at the response to a small external field perturbation $h^{\prime}$, after cooling the system to the desired low temperature. In contrast, the equilibrium susceptibility $\chi_{eq}$ is approximated by the field cooled susceptibility, which is measured by applying the small field perturbation already while cooling the system below the spin glass transition temperature. In this case, the system has time to explore and select the most appropriate state while cooling in presence of the external perturbation. Experimental results of the two different spin glass susceptibilities below the critical temperature are illustrated for Cu(Mn13.5%) in Fig. 8 and compared with theoretical predictions. 90, 78
@@
Figure 8: Comparison of experimental (top) and model (bottom) susceptibilities with respect to temperature $T$. The linear response susceptibility ($\chi_{LR}$, lower curve) can be measured experimentally by applying an external field perturbation after the cooling of the material (Cu(Mn13.5%)) below the critical spin glass temperature. In contrast, the equilibrium susceptibility ($\chi_{eq}$, upper curve) can be approximated experimentally by applying the magnetic field perturbation before the cooling. 90, 78 Above the critical temperature the material enters the paramagnetic stable phase, with only one susceptibility, i.e., the linear response accesses equilibrium properties.
@@
In spin glass theory, e.g., the SK model, there is a clear distinction between replica symmetry breaking and hysteresis effects: 78 Hysteresis is commonly attributed to defects that are localized in space and induce a finite free energy barrier and thus finite lifetime of meta-stable states. Thus, in hysteresis, the two susceptibilities coincide after waiting sufficient long time. In contrast, the non-local barriers in spin glasses imply the re-arrangements in arbitrary large regions of the system, which can even diverge in the thermodynamic limit. Therefore, the different susceptibilities will not agree, provided that the externally applied field perturbation is sufficiently small.
@@
Whether such a clear distinction between hysteresis and replica symmetry breaking also holds in a polaritonic setup is unknown and requires more experimental and theoretical work to demonstrate that polaritonic chemistry behaves as a spin glass system. 91 As we can deduce from the previous discussions, the quasi-static spin-glass picture of the SK model is incomplete for our polaritonic setup, since the dressed electronic structure of an ensemble of molecules under VSC will be (periodically) driven by the dynamics of the nuclear and displacement field coordinates. Therefore, it is likely that time-dependent external field perturbations are required to probe the coexistence of different cavity-induced linear (or higher-order) susceptibilities (see discussion in Sec. 4.2.2). In particular, we already see that Eq. (44) should explicitly depend on time, since the Edwards-Anderson order parameter of the cH equations directly depends on $(R(t),q_{\beta}(t))$ and thus $q_{EA}(t)$. If this also implies the coexistence of different (possibly dynamic) susceptibilities under VSC remains a non-trivial open question. A further hint at interesting effects is the fact that we have now a highly degenerate ground state which makes response calculations intricate. However, investigating dynamic susceptibilities could provide a promising route to verify and characterize the proposed polarization-glass phase in a cavity.
4.2.2 Off-equilibrium: Time-correlations and cavity-induced feedback effects
Even more interesting than the equilibrium properties of spin glasses are their associated off-equilibrium phenomena, which are observable for different materials. In order to measure the generalized susceptibility, it is common practice to rely on the fluctuation-dissipation theorem, which relates the response of the system to an external perturbation (weak off-equilibrium) to its equilibrium properties (fluctuations/correlations). The fluctuations of glassy systems can be characterized by the time-correlations of the magnetizations (or polarizations respectively),
$C(t,t_{w})=\frac{1}{N}\sum_{i}^{N}\langle\sigma_{i}(t_{w})\sigma_{i}(t_{w}+t)\rangle\Rightarrow\frac{1}{N}\sum_{i}^{N}\frac{\langle\Delta\mu_{i}(t_{w})\Delta\mu_{i}(t_{w}+t)\rangle}{\langle\Delta\mu_{i}(t_{w})\Delta\mu_{i}(t_{w})\rangle}.$ (46)
Comparing the magnetization time-correlations of a (quasi-static) spin glass with the electronic correlations of a polarization glass reveals similarities (see Fig. 9) but again highlights some fundamental physical differences. In a spin glass, the magnetization correlations decay monotonically, but extremely slowly even on a logarithmic time-scale. In addition, the random process of $\sigma_{i}(t)$ cannot be considered a wide-sense stationary stochastic process, which means the correlation does not only depend on the time-difference $t$, but also on the waiting-time $t_{w}$ elapsed since entering the spin-glass phase. In other words, time-reversal symmetry is explicitly broken in a spin glass. The explicit waiting-time dependency will give rise to specific aging effects that are discussed in Sec. 4.2.3. In contrast, the polarization glass correlations reveal different oscillatory regimes, introduced by the time-dependent $(R(t),q_{\beta}(t))$-parameters. MD simulations with $N=36$ Shin-Metiu molecules reveal the following correlation features within an adiabatic regime:
  1. 1.
    Long-lived periodic (dynamic) correlation patterns emerge.
  2. 2.
    Fast oscillations (vibrational and cavity time scales) that are symmetrically distributed around $C(t)=0$ (light-blue), i.e., the existence of anti-correlated states.
  3. 3.
    The amplitude of the fast oscillations is modulated periodically on an order of magnitude smaller frequency scale (blue), with even coarser intensity modulations on a pico-second time scale.
  4. 4.
    The long-lived correlations are introduced by the harmonic nature of the strongly coupled cavity mode $q_{\beta}$ and their overall magnitude depends on the losses of the cavity (a Langevin-thermostat is coupled to $q_{\beta}$, see Ref. 20 for implementation details). In particular, a more Brownian-like motion of $q_{\beta}$ effectively prevents that long-lived correlations emerge. The Brownian dynamics of $q_{\beta}$ can be interpreted as the weak coupling to a continuum of modes or thermal bath instead.
Investigating thoroughly the long-lived correlation dynamics in a polarization glass is already a computationally very demanding task. Particularly, the deviations of $\Delta\mu_{i}(t)$ from a wide-sense stationary stochastic process are non-trivial to detect and thus remain an open research question. In particular, significant deviations may potentially only occur when increasing the collective coupling strength beyond the onset of the polarization glass instability. The breakdown of the adiabatic picture for the dressed electronic structure will increase the computational complexity even further in this regime, which may require the development of novel computational algorithms. However, the strongly oscillatory nature of the correlations within the adiabatic regime may also have immediate implications for the development of a dynamic susceptibility picture, as previously discussed. When averaging the (immediate) linear response susceptibilities over many oscillatory cycles, little to no effect (deviation from the mean) is expected. This suggests that higher-order response functions will likely become relevant observables to unravel the inner workings of a polarization glass and thus VSC in general. Overall, characterizing correlations, (non-linear) responses as well as non-adiabatic effects will have important implications for the off-equilibrium properties of a cavity-induced polarization glass. Particularly, non-adiabatic effects are known to essentially determine the physics of the SK model, as we detail in the following.
Figure 9: Comparison of time-correaltions in a spin and polarization glass: Top: Monotonically decaying time correlation functions $C(t,t_{w})$ of an Ising spin glass (taken from Ref. 78 based on data from Ref. 92). The dependency on the waiting time $t_{w}$ indicates the aging behaviour. Bottom: Oscillating polarization correlations of the Shin-Metiu setup under VSC with $N=36$ molecules. When calculating $C(t)$ it was implicitly assumed that $\Delta\mu_{i}(t)$ is a wide-sense stationary stochastic process within the adiabatic cH picture. The correlations oscillate fast (light-blue), on the time-scale of the cavity (vibrations), where the bold lines visualize the enveloping function. The blue line corresponds to the strong coupling to one lossy cavity mode ($q_{\beta}$ weakly coupled to a Langevin bath). In contrast, the coupling to a Brownian mode is shown in red, which can be interpreted as the weak coupling to many cavity modes (thermal bath) instead.
@@ 4.2.2 Off-equilibrium: Time-correlations and cavity-induced feedback effects
4.2.3 Off-equilibrium: Breakdown of the fluctuation-dissipation theorem and aging effects
@@
Let us next make the following assumptions: At time $t_{0}=0$ the system is suddenly cooled below the critical glass temperature $T<T_{g}$, which triggers the phase transition into a spin glass. After a waiting time $t_{w}$, a constant external field perturbation $h^{\prime}$ is applied. Eventually, the response of the system $S(t_{w},t)$ is evaluated at time $t+t_{w}\geq t_{w}$. Under these conditions, the fluctuation-dissipation relations connect 93, 94, 78 the time correlation function of the total magnetization $C(t,t_{w})$, defined in Eq. (46), to the average relaxation function per spin, which is defined by 78
@@
Eq. (47) describes the response of the magnetization at time $t_{w}+t$ if the external field was added at time $t_{w}$. However, the standard fluctuation-dissipation theorem is only applicable if the system obeys the detailed balance condition, 95 which can be violated in a glassy system on longer time-scales. To account for this aspect, modified fluctuation dissipation relations have been proposed that hold for weak off-equilibrium. 96, 97, 75 In more detail, two different off-equilibrium regimes are distinguished for (spin) glasses, for which different fluctuation-dissipation relations hold: 94
In the stationary correlation regime of a spin glass, the correlations are assumed to solely depend on $t$, but not on the waiting time $t_{w}$, i.e, $C(t,t_{w})=C_{s}(t)$. Typically this approximation is reasonable only for relatively small $t\approx 0$ in spin glasses, which implies high correlations $C_{s}(t)\approx 1$. Having correlations close to unity, the standard (thermal equilibrium) fluctuation-dissipation relation are applicable, which yields 92, 98
@@
In the aging regime, where the correlations are no longer stationary, modified fluctuation-dissipation relations of the following form were suggested 75
@@
In other words, the deviations of the fluctuation-dissipation theorem that are caused by aging effects can be related to equilibrium properties given by $P(\mathscr{q})$. An illustration of the two different off-equilibrium regimes with their relation to equilibrium properties is given in Fig. 10(a) for the SK model in comparison with standard hysteresis effects, i.e., visualizing the difference between replica symmetry breaking and hysteresis. Notice the modifications of the fluctuation-dissipation relations can be re-interpreted in terms of an effective temperature 100, 97
@@
which indicates a heating or excess of thermal fluctuations since $0\leq dS/dC\leq 1$ according to the probability interpretation of Eq. (50). The two different off-equilibrium regimes, i.e., the emergence of aging effects, seem to be a generic feature of glassy systems. 75 In Fig. 10(b) experimentally recorded fluctuation-dissipation relations are shown for CdCr1.7In0.3S4 with respect to different finite waiting times. 102 The extrapolation to infinite waiting times allows the connection to the theoretical spin glass models (i.e., the connecting to equilibrium properties).
Including the parametric feedback of the nuclei on the dressed electronic structure modifies the fluctuation-dissipation picture of a spin glass significantly. In particular, the emergence of periodic oscillations implies that the thermal equilibrium fluctuation-dissipation relation, given in Eq. (48), will no longer be applicable in the stationary correlation picture. This is not so surprising since the electronic subsystem is strongly correlated with the other subsystems (not decoupled, i.e. strongly hybridised). Therefore, Eq. (48) will remain valid solely for very short time-scales ($C_{s}(t)\approx 1$), or if the cavity is too lossy/too weakly coupled to the system to induce any chemical changes on the molecular ensemble (see suppressed correlations visualized by the red line in Fig. 9). Consequently, different fluctuation-dissipation relations are expected even in the stationary correlation regime under VSC. Similarly, by increasing the collective coupling strength further, the breakdown of the adiabatic approximation suggest the emergence of a dynamic aging regime under VSC. In analogy to the stationary case, we expect that the dynamic feedback effects require modifications of the fluctuation-dissipation relation in Eq. (49) to be captured.
4.2.4 Polaritonic resonance picture revised

### v2

@@ 4.1.3 Equilibrium Susceptibilities
The hallmark of replica symmetry breaking in spin glass theory can be attributed to the emergence of two different static equilibrium susceptibilities, 118 which are observed in experiments 130 and in the SK model. 108 The two extreme cases describe the response of the system subject to a small external field perturbation. In the so-called zero-field cooled case, the system remains inside a given state while changing the magnetization (electric, for the polaritonic setup) field, with a corresponding linear response susceptibility $\chi_{LR}$. In contrast, the (true) thermodynamic equilibrium susceptibility $\chi_{eq}$ describes the situation, where the spin glass is allowed to relax to the thermodynamically most favored state in presence of a weak external field perturbation.
@@
Experimentally, the static linear-response susceptibility $\chi_{LR}$ can be measured by looking at the response to a small external field perturbation $h^{\prime}$, after cooling the system to the desired low temperature. In contrast, the equilibrium susceptibility $\chi_{eq}$ is approximated by the field cooled susceptibility, which is measured by applying the small field perturbation already while cooling the system below the spin glass transition temperature. In this case, the system has time to explore and select the most appropriate state while cooling in presence of the external perturbation. Experimental results of the two different spin glass susceptibilities below the critical temperature are illustrated for Cu(Mn13.5%) in Fig. 7130, 118
@@
Figure 7: Experimental susceptibilities with respect to temperature $T$. The linear response susceptibility ($\chi_{LR}$, lower curve) can be measured experimentally by applying an external field perturbation after the cooling of the material (Cu(Mn13.5%)) below the critical spin glass temperature. In contrast, the equilibrium susceptibility ($\chi_{eq}$, upper curve) can be approximated experimentally by applying the magnetic field perturbation before the cooling. 130, 118 Above the critical temperature the material enters the paramagnetic stable phase, with only one susceptibility, i.e., the linear-response accesses equilibrium properties.
@@
In spin glass theory, e.g., the SK model, there is a clear distinction between replica symmetry breaking and hysteresis effects: 118 Hysteresis is commonly attributed to defects that are localized in space and induce a finite free-energy barrier and thus finite lifetime of meta-stable states. Thus, in hysteresis, the two susceptibilities coincide after waiting sufficiently long time. In contrast, the non-local barriers in spin glasses imply the re-arrangements in arbitrary large regions of the system, which can even diverge in the thermodynamic limit. Therefore, the different susceptibilities will not agree, provided that the externally applied field perturbation is sufficiently small.
@@
A clear distinction between hysteresis and a spin glass is not so trivial in a polaritonic setup, as we will see in Sec. 4.2, and thus requires more experimental and theoretical work. 131 Moreover, as we can deduce from the previous discussions, the quasi-static spin glass picture of the SK model is incomplete for our polaritonic setup, since the dressed electronic-structure of an ensemble of molecules under VSC will be (periodically) driven by the dynamics of the nuclear and displacement field coordinates. Therefore, it is likely that time-dependent external field perturbations are required to probe the coexistence of different cavity-induced linear (or higher-order) susceptibilities. In particular, we already see that Eq. (50) should explicitly depend on time, since $J_{ij}$ and thus the Edwards-Anderson order parameter depend on $(R(t),q_{\beta}(t))$ and thus $q_{EA}(t)$. If this also implies the coexistence of different (possibly dynamic) susceptibilities under VSC remains a non-trivial open question. A further hint at interesting effects is a highly degenerate ground state, which makes response calculations intricate. However, investigating dynamic susceptibilities could provide a promising route to verify and characterize the proposed polarization-glass phase in a cavity.
@@
4.1.4 Off-equilibrium spin glass: Breakdown of the fluctuation-dissipation theorem and aging effects
Even more interesting than the equilibrium properties of spin glasses are their associated off-equilibrium phenomena, which are observable for different materials. In order to measure the generalized susceptibility, it is common practice to rely on the fluctuation-dissipation theorem, which relates the response of the system to an external perturbation (weak off-equilibrium) to its equilibrium properties (fluctuations/correlations). The fluctuations of glassy systems can be characterized by the time-correlations of the magnetizations (or delocalized/intermolecular excitations and their induced polarizations, respectively, for the polaritonic setup),
$C(t,t_{w})=\frac{1}{N}\sum_{i}^{N}\langle\sigma_{i}(t_{w})\sigma_{i}(t_{w}+t)\rangle$ (52)
In a spin glass, the magnetization correlations decay monotonically, but extremely slowly even on a logarithmic time-scale. In addition, the random process of $\sigma_{i}(t)$ cannot be considered a wide-sense stationary stochastic process, which means the correlation does not only depend on the time-difference $t$, but also on the waiting-time $t_{w}$ elapsed since entering the spin glass phase. In other words, time-reversal symmetry is explicitly broken in a spin glass. The explicit waiting-time dependency will give rise to specific aging effects that are discussed in in the following:
Figure 8: Comparison of aging behavior for two spin glass models with (left) and without (right) replica symmetry breaking. Stationary correlation regime is marked in light blue and the aging regime in light red. Left: Monotonically decaying time correlation functions $C(t,t_{w})$ of an Ising spin glass (taken from Ref. 118 based on data from Ref. 132). The dependency on the waiting time $t_{w}$ indicates the aging behavior, i.e., the deviation from $C_{\rm eq}(t)$. Right: Time-correlation functions of the Spherical Sherrington Kirkpatrick (SSK) model. Notice the plateau region of the autocorrelation function at $C=q_{\rm EA}$, which depends asymptotically on the waiting time, thus the correlations may not necessarily decay to zero. Data taken from Ref. 133.
@@ 4.1.4 Off-equilibrium spin glass: Breakdown of the fluctuation-dissipation theorem and aging effects
For simplicity, let us make the following assumptions: At time $t_{0}=0$ the system is suddenly cooled below the critical glass temperature $T<T_{g}$, which triggers the phase transition into a spin glass. After a waiting time $t_{w}$, a constant external field perturbation $h^{\prime}$ is applied. Eventually, the response of the system $S(t_{w},t)$ is evaluated at time $t+t_{w}\geq t_{w}$. Under these conditions, the fluctuation-dissipation relations connect 134, 135, 118 the time correlation function of the total magnetization $C(t,t_{w})$, defined in Eq. (52), to the average relaxation function per spin, which is defined by 118
@@
Eq. (53) describes the response of the magnetization at time $t_{w}+t$ if the external field was added at time $t_{w}$. However, the standard fluctuation-dissipation theorem is only applicable if the system obeys the detailed balance condition, 136 which can be violated in a glassy system on longer time-scales (e.g., by spontaneous replica symmetry breaking). To account for this aspect, modified fluctuation dissipation relations have been proposed that hold for weak off-equilibrium . 137, 138, 117 In more detail, two different off-equilibrium regimes are distinguished for (spin) glasses, for which different fluctuation-dissipation relations hold: 135
In the stationary correlation regime of a spin glass, the correlations are assumed to solely depend on $t$, but not on the waiting time $t_{w}$, i.e, $C(t,t_{w})=C_{s}(t)$. Typically this approximation is reasonable only for relatively small $t\approx 0$ in spin glasses, which implies high correlations $C_{s}(t)\approx 1$. Having correlations close to unity, the standard (thermal equilibrium) fluctuation-dissipation relation are applicable, which yields 132, 139
@@
In the aging regime, where the correlations are no longer stationary, modified fluctuation-dissipation relations of the following form were suggested 117
@@
In other words, the deviations of the fluctuation-dissipation theorem that are caused by aging effects can be related to equilibrium properties given by $P(\mathscr{q})$. An illustration of the two different off-equilibrium regimes with their relation to equilibrium properties is given in Fig. 9 for the SK model in comparison with standard hysteresis effects, i.e., visualizing the difference between replica symmetry breaking and hysteresis. Note that the modifications of the fluctuation-dissipation relations can be re-interpreted in terms of an effective temperature 141, 138
@@
which indicates a heating or excess of thermal fluctuations since $0\leq dS/dC\leq 1$ according to the probability interpretation of Eq. (56). The two different off-equilibrium regimes, i.e., the emergence of aging effects, seem to be a generic feature of glassy systems. 117 In Fig. 10 experimentally recorded fluctuation-dissipation relations are shown for CdCr1.7In0.3S4 with respect to different finite waiting times. 143
4.2 Spherical Sherrington-Kirkpatrick (SSK) model
The spherical Sherrington-Kirkpatrick (SSK) model (or spherical 2-spin glass) closely resembles the SK model introduced in Eq. (34). However, the SSK model possesses continuous spin variables $s_{i}$, which additionally obey the following “spherical” constraint $\sum_{i}s_{i}^{2}=N_{J}$ that ensures a finite ground state energy. The SSK model was introduced by Kosterlitz, Thouless, and Jones in Ref. 109 as
$H_{SSK}(\boldsymbol{s})=-\sum_{i<j}^{N_{J}}s_{i}s_{j}J_{ij},\ J\sim\mathscr{N}(J_{0}/N_{J},\tilde{J}^{2}/N_{J}).$ (58)
For simplicity, we assume a symmetric random matrix, i.e, $J_{ij}=J_{ji}$ in the following, as it is the case for the DSE correlation energy. In contrast to the SK model, the SSK model is easier to analyze144. In particular, a simple analytic solution for the free-energy was found using random-matrix theory.109 In the absence of an external magnetic field, the expected free energy per site (averaged over different realizations of $J_{ij}$) is given as109
$F_{T,\tilde{J}}=\lim_{N_{J}\rightarrow\infty}F_{N_{J},T,\tilde{J}}=\begin{cases}-\frac{\tilde{J}^{2}}{4k_{B}T}-\frac{1}{2}k_{B}T(1+\ln(2))&\text{if }T\geq T_{c},\ J_{0}=0\\ \frac{1}{2}k_{B}T\ln\big{(}k_{B}T/(2\tilde{J})\big{)}-\tilde{J}+\frac{1}{4}k_{B}T&\text{if }T<T_{c},\ J_{0}=0,\end{cases}$ (59)
in the thermodynamic limit. The Edwards-Anderson order parameter is given by $q_{EA}^{\rm SSK}=1-T/T_{c}$. It reveals a spin glass phase for temperatures below the critical temperature $T_{c}=\tilde{J}/k_{B}$. From a static point of view the model is relatively simple, since the SSK model does not break replica symmetry.133 In particular, its energy has only two minima,133 which suggests the absence of frustration and thus trivial dynamics. However, it turns out that the energy-fluctuations 144 and the dynamics 133 of the SSK model are modified non-trivially. Indeed, significant aging effects appear and for almost any initial condition, the glassy system is out of equilibrium, and the evolution does not (!) lead to equilibrium. 133.
In more detail, the order (decaying with respect to the number of spins) and the distribution of the free-energy fluctuations per spin among the ensemble changes as follows,
$F_{T,1}-F_{N_{J},T,1}\propto\begin{cases}\frac{1}{N_{J}}\mathscr{N}&\text{if }T\geq T_{c},\ J_{0}=0\\ \frac{1}{N_{J}^{2/3}}TW_{1}&\text{if }T<T_{c},\ J_{0}=0,\end{cases}$ (60)
[…]
For the SSK model, analytic solutions can be found for the two-time spin correlation function defined in Eq. (52) in the spin glass phase. 133 For this purpose, Cugliandolo and Dean propagated the spin dynamics (defined via Langevin equations) for different initial conditions. Surprisingly, for almost any initial condition non-equilibrium dynamics occurs! Only for very specific initial conditions, with “a macroscopic condensation” determined by the eigenvalue of the random matrix $J_{ij}$, equilibrium dynamics is achieved. 133 In the following, we focus on the spin correlations of the generic non-equilibrium setup, for which we recover the stationary correlation regime for relatively small $t$ and the aging regime, where the correlations explicitly depend on the waiting time $t_{w}$ (see Fig. 8). The stationary correlation regime
$C(t,t_{w})\sim C_{s}(t),\ t\ll t_{w}.$ (61)
is characterized by a fast initial decay from $C=1$ to $C=q_{EA}$ for $t\ll t_{w}$, where the usual fluctuation dissipation relations hold according to Eq. (54). The aging regime can be subdivided into two further sub-regimes as follows,
$C(t,t_{w})\sim\begin{cases}q_{EA}^{\rm SSK}=1-T/T_{c}&\text{if }t\approx t_{w},\\ \rm{slowly\ decaying\ to\ 0}&\text{if }t\gg t_{w}.\end{cases}$ (62)
One of the striking features of the SSK model is that the plateau region of the autocorrelation function in Eq. (62), depends asymptotically on the waiting time, i.e., 133
$\lim_{t_{w}\rightarrow\infty}C(t,t_{w})=q_{EA},\ \forall t\ \rm{finite}.$ (63)
Furthermore, the decay of the correlation for fixed $t\gg t_{w}$ is extremely slow (notice the logarithmic time-scale in Fig. 8).
However, despite above long-lived correlations, aging effects in the magnetization are weak, i.e., the relaxation of an externally induced magnetization $S$ defined in Eq. (53) decays fast. As a consequence, one can show that the modified fluctuation-dissipation relation are sort of trivial and one finds 133
$\frac{dS}{dC}=X(C)\sim\begin{cases}-1&\text{if}\ C(t,t_{w})\sim C_{s}(t),\\ 0&\text{otherwise, }\end{cases}$ (64)
as visualized in Fig. 9. The condition $\frac{dS}{dC}=0$ suggest a divergent effective temperature $\tau\rightarrow\infty$, i.e., a strong excess of thermal fluctuations in the aging regime, which we interpret as a sign of the persistent non-equilibrium dynamics of the SSK model. At the same time, the linear-response susceptibility matches the equilibrium susceptibility $\chi_{LR}^{\rm SSK}=\chi_{eq}^{\rm SSK}$, similar to standard hysteresis effects. Thus applying a small static external field perturbation does not probe the complex correlations of the SSK model. As previously mentioned, applying time-dependent external fields may be an interesting option for VSC. However, at the moment it remains unclear if they could probe dynamic correlation effects in a SSK setup.
5 Interpreting polaritonic chemistry with the SSK model
The similarities between the SSK model, defined in Eq. (58), and the cavity-mediated electron correlations in Eq. (33), are striking. In the following section, we derive different theoretical consequences of the potential spin glass nature of cavity-mediated electron correlations, and present a concise picture of its consequences for chemistry under VSC. Eventually, we briefly connect our theoretical picture to recent experimental result.
5.1 SSK electron correlations
The aforementioned similarity suggests the following assignment for the total free energy of the intermolecular DSE correlations
$F_{\rm corr}^{\rm DSE}\sim N_{J}F_{T,\tilde{J}}.$ (65)
This implies that we explicitly impose normally-distributed random interactions for $J_{ij}$. Realistic molecular ensembles in a cavity will certainly deviate considerably, and we expect rather heavy-tailed distributions to appear in nature, i.e., relatively few degrees of freedom contribute strongly to $J_{ij}$, whereas the others are only marginally (rare-event driven). Determining realistic distributions is non-trivial and requires considerable future research efforts. Clearly, the spin glass properties will be affected by the distribution. Nevertheless, the overall picture developed in Sec. 4 should still apply, since the generic features of spin glasses remain preserved qualitatively across many known models in the literature. 117 As an immediate consequence of collective strong coupling, thermal effects must be included when solving the dressed electronic problem, even-though the thermal energy scale can be orders of magnitude smaller than the free-space excitation of the electronic-structure of a single molecule. This can be rationalized by the fact that the novel collective electronic excitations can be on a much lower energy scale, similar to solid-state systems. In particular, we expect a cavity-induced spin glass phase transition to occur if all prerequisites, discussed in Sec (3), are met by the polaritonic system. Indeed, Eq. (60) suggest differently distributed (and differently scaling) fluctuations of the electronic correlations for $T<T_{c}$. Notice, while formally the changed fluctuation is localized per spin, this does not imply that the corresponding intermolecular DSE correlations is localized in space. Indeed, we expect that rather de-localized orbitals are affected, which contribute to the long-range intermolecular dispersion effects. Those (rare events) are expected to allow for significant overlap integrals, represented by the tails of the $J_{ij}$-distribution. Nevertheless, the proposed spin glass phase transition provides a theoretical mechanism, which could explain, why under many circumstances no chemical changes are observed, i.e., $T>T_{c}$. However, under specific but non-trivial conditions, $T<T_{c}$ can be reached, which implies that rare events and fluctuations scale differently. These changes can become chemically relevant under collective strong-coupling conditions.
Apart from the modified fluctuations in a static picture, the SSK mapping suggests unique dynamic features. In particular, the emergence of long-lived time-correlations and aging effects (see Figs. 8 & 9) effectively prevents the DSE electron correlation dynamics to (ever) reach thermal equilibrium even for long waiting times at $T<T_{c}$. This suggests a cavity-induced time-reversal symmetry breaking. This could also open novel pathways for the sub-field of chiral polaritonics. 146 There the common aim is to reach cavity-induced enantioselctivity by explicitly parity-violating cavity polarizations. 147, 148, 149, 150, 151, 152, 153, 154 However, from a fundamental theoretical aspect similar effects may also be reached by spontaneous symmetry breaking instead 155, 156 and/or degenerate states.146 At this point it is important to mention that overall our dynamic picture remains incomplete. In particular, in a polaritonic setup, the entire physical system will evolve, i.e., the nuclear and displacement field parameters $R,q_{\beta}$ will explicitly depend on time and thus $J_{ij}\mapsto J_{ij}(t)$, which is not considered by the discussions in Sec. 4. In that sense, our dynamic picture of the electrons can only be considered as the limiting case of infinitely-slow dynamics of the external parameters (quasi-static). How the different involved time-scales favor or suppress the emergence of non-equilibrium aging effects remains an open question. However, the emergence of non-equilibrium electron dynamics could have fascinating consequences on the thermal equilibrium features of the entire ensemble, as we would like to briefly argue in the following.
5.2 Non-canonical nuclear dynamics / stochastic resonances
Returning to our initial Born-Oppenheimer partitioning, we notice that approximating the ”ground-state” dynamics of the nuclei and displacement field classically (see Eq. (5)) becomes considerably more complex. Even when ignoring non-adiabatic couplings for the highly degenerate polaritonic ground-state, i.e., sustaining a classical picture, the non-equilibrium electron dynamics in the aging regime implies that the electronic force contributions become explicitly time-dependent within the spin glass phase. Therefore,
$\mathscr{E}\big{(}\boldsymbol{R}(t),q_{\beta}(t)\big{)}\overset{T<T_{c}}{\mapsto}\mathscr{E}\big{(}\boldsymbol{R}(t),q_{\beta}(t),t\big{)}.$ (66)